Three 101.0-g ice cubes initially at 0°C are added to 0.820 kg of water initially at 19.5°C in an insulated container. (a) What is the equilibrium temperature of the system? °C (b) What is the mass of unmelted ice, if any, when the system is at equilibrium? kg
Answer #1

Here,

let the equilibrium temperature of the system is T

heat needed to melt ice = 3 * 101 * 334 = 101202 J

heat released when the water reaches 0 degree C = 0.820 * 4186 * (19.5 - 0)

heat released when the water reaches 0 degree C = 66934 J

hence , there is not enough heat to melt the ice

the final temperature will be 0 degree C

part B)

let the unmelted ice is m

heat released = (3 * 101 - m) * 334

66934 = (303 - m) * 334

m = 102.6 gm

the mass of ice remaining is 102.6 gm